Find an approximate value for the rate at which the moon orbits the earth. Assume that the moon's orbit is circular.: Find an approximate value for the rate at which the moon orbits the earth. Assume that the moon's orbit is circular.
Answer by MathLover1(1157)

an_approximate_value_for_the_rate =2r*pi/ the_amount_of_days
since the distance between the earth and moon is 382000000 and the amount of days it takes the moon to orbit earth is 27.3, we will have
an_approximate_value_for_the_rate =2*(3.14) * 382000000/27.3 = 2398960000 /27.3
an_approximate_value_for_the_rate =87873992.67 (meters/day)

But, the moon's orbit is elliptical, not circular. So, above result is close. It should be a larger than that because of the elliptical orbit as opposed to circular.
One way you could look at it is that the mean orbital speed of the moon is 1023 m/s.
There are 86400 s in a day, therefore
an_approximate_value_for_the_rate = (86400)(1023)=88387200 (meters/day).